A woman had $1270 in her checking account. She deposited $180 and wrote checks for $25 and $110. She then wrote a check for $45 and deposited $150. How much did she finally have in her account? Write as an addition expression using integers.
Answer by jorel555(1290)

Total deposits:1270+180+150=1600
Total withdrawals:25+110+45=180
Ending balance: 1600-180=$1420. ☺☺☺☺